# VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition

### v3

@@ 3.4 Character and String Modeling
Existing studies mainly use feature sequences with height $1$ to represent image text. Each feature corresponds to a small local region. However, these segments are usually noisy, especially for irregular text, failing to describe the characters of an image well. In contrast, the Vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds. This suggests that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual dependencies of strings.
@@
More specifically, image text requires features to be extracted from two perspectives. The first involves local character features, including elements like strokes and fonts, representing internal associations of a character and morphological features. The second is context relations, such as correlations between different characters or between text and non-text components. Therefore, we design two classes of attention-mixing blocks with different receptive fields to sense relevance. Mathematically speaking, for a sequence of character components $B_{i-1}B_{i}$ in the previous stage, it is first reshaped as a sequence of features. Upon the input of the attention mixing block, Layer Normalization (LN) and Position Embedding (PE) are first applied. Subsequently, the self-attention mechanism is employed for dependency modeling. Finally, LN and Feed Forward Network (FFN) are successively applied for feature fusion, with the residual connection applied.
Positional Embedding.
@@
Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back. Typical positional Embedding mechanisms include absolute positional encoding (APE) Vaswani et al. (2017), relative positional encoding (RPE) Liu et al. (2021), and CPE. APE and RPE are designed for a specific input size and are not friendly to varying input resolutions. CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions. Then the generated positional encoding will be added to the input feature. Locally-enhanced Positional Encoding (LePE) Dong et al. (2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
Local Attention.
@@
The goal of Local Attention is to encode character morphological features and establish correlations within character components, thereby simulating stroke features. Local Attention needs to consider the neighborhood of character components on the two-dimensional coordinate system. Drawing on the design ideas of MaSA and CSWin Self-Attention, we designed two efficient self-attention mechanisms that explicitly model local features on a two-dimensional coordinate system. MaSA is shown in Figure 3LABEL:sub@subfig:dmsa. The spatial attenuation matrix of the image means that character features are highlighted globally and background features are weakened. CSWin Self-Attention is shown in Figure 3LABEL:sub@subfig:cswin. This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel. For image text input, considering that the image resolution is larger than the Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing, and the length and width resolution needs to be guaranteed. Divisible by $SW$, we empirically set $SW$ for stages using CSWin Self-Attention to $1$ or $2$ by default.
Global Attention.
@@
As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual dependencies between all character components. Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters. In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components. In addition, to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information. The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
As mentioned above, the hybrid blocks in VIPTR are designed to extract distinct yet complementary features. In this way, blocks are initially processed locally before being addressed globally. At each stage, Local and Global Attention are recurrently applied multiple times for comprehensive feature extraction. Effect of the two kinds of blocks Permutations can be seen in ablation experiments.
3.5 Subsampling
To save computational costs and extract significant features layer by layer, we added a down-sampling operation Du et al. (2022) on the height dimension after the first and third attention mixing stages of VIPTR. Using the output of the features from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively. We then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$. The Subsampling block halves the height while maintaining a constant width. It not only reduces the computational cost, but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension. We also increase the channel dimension $C_{i}$ to compensate for the information loss.
3.6 Combining and Prediction
@@
In the final stage, we use combinatorial operations Du et al. (2022) to integrate sequence features. It first uses AdaptiveAvgPool to compress the height dimension to $1$, followed by fully connected layers, Hardswish activation, and dropout. By doing so, the character components are further compressed into feature sequences, where each element is represented by a feature of length $C_{3}$. Using AdaptiveAvgPool avoids applying convolution to very small embeddings of one dimension, such as height $2$, while being able to apply a variety of initial input heights.
@@
3.7 Vision Permutation and Architecture Variants
Firstly, as shown in Figure 2, in the 2-3 stages of VIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called VIPTRv1 and VIPTRv2. Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, and number of mixing blocks. By varying them, VIPTR architectures with different numbers of parameters can be obtained. In summary, we constructed four typical architectures, namely VIPTRv1-T (Tiny), VIPTRv2-T, VIPTRv1-L (Large), and VIPTRv2-B (Base). Its detailed configuration is shown in Table 2.
@@ 3.7 Vision Permutation and Architecture Variants
CRNN (None) 75.3 56.7 0.69
SVTR-T 77.9 62.5 0.29
$[L_{1}][L_{1}G_{1}][G_{1}]$ 79.1 65 0.269
$[L_{1}][L_{1}G_{2}][G_{1}]$ 79.1 65.2 0.258
$[L_{2}][L_{2}G_{1}][G_{1}]$ 79 65.1 0.282
$[L_{2}][L_{2}G_{2}][G_{1}]$ 79.3 65.3 0.27
$[L_{1}][L_{2}G_{1}][G_{1}]$ 78.8 64.9 0.27
$[L_{2}][L_{1}G_{1}][G_{1}]$ 78.9 65.1 0.27
$[L_{1}][L_{2}G_{2}][G_{1}]$ 78.7 65 0.258
$[L_{2}][L_{1}G_{2}][G_{1}]$ 78.8 64.8 0.258
$[L_{1}][L_{1}//G_{1}][G_{1}]$ 75.5 60.5 0.195
$[L_{1}][L_{1}//G_{2}][G_{1}]$ 75.4 60.5 0.192
$[L_{2}][L_{2}//G_{1}][G_{1}]$ 75.6 60.4 0.197
$[L_{2}][L_{2}//G_{2}][G_{1}]$ 75.8 61 0.194
$[L_{1}][L_{2}//G_{1}][G_{1}]$ 75 59.9 0.196
$[L_{2}][L_{1}//G_{1}][G_{1}]$ 75.2 59.7 0.196
$[L_{1}][L_{2}//G_{2}][G_{1}]$ 74.9 60.1 0.193
$[L_{2}][L_{1}//G_{2}][G_{1}]$ 75.2 60.3 0.193
Table 1: Ablation study on Attention Blocks’ permutations.
VIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 4.0 0.26
VIPTRv1-L $[192,256,512,384]$ $[3,7,2,9]$ $[6,8,8,16]$ $[L_{2}][L_{2}G_{2}][G_{1}]$ 37.7 2.31
VIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.2 0.19
VIPTRv2-B $[128,256,384,256]$ $[3,6,6,9]$ $[4,8,8,12]$ $[L_{2}][L_{2}//G_{2}][G_{1}]$ 20.2 1.18
Table 2: VIPTR architecture variants (w/o counting the rectification module and CTC decoder).
4 Experiments and Discussion
@@
Our English STR models are trained on two synthetic datasets, MJSynth (MJ) Jaderberg et al. (2014, 2016) and SynthText (ST) Gupta et al. (2016) trainsets (72k+55k), along with real-world data from COCO-Text Gomez et al. (2017). Testing is conducted on six public benchmarks: ICDAR 2013 (IC13) Karatzas et al. (2013) testset with 857 regular images, Street View Text (SVT) Wang et al. (2011) testset with 647 images from Google Street View, IIIT5K-3000 Mishra et al. (2012) sourced from websites, ICDAR 2015 (IC15) Karatzas et al. (2015) comprising 1811 images captured via Google Glasses, Street View Text Perspective (SVTP) Phan et al. (2013) with 639 images also from Google Street View, and CUTE80 (CUTE), which focuses on curved text recognition, with 288 images.
For the Chinese recognition task, we employ the Chinese Scene and Document Dataset Yu et al. (2021). The former includes a scene text dataset with 509,164 training, 63,645 validation, and 63,646 test images, featuring 5,880 unique characters. The document text dataset comprises 400,000 training, 50,000 validation, and 50,000 test images, with 4,865 distinct characters.
Implementations
The recommended training and testing configurations for VIPTR are listed in Table 3. When Training English models, data augmentation like stretch, perspective distortion, blur, and Gaussian noise, are randomly performed with probability $0.4$. Then, It was not used when training the Chinese model. All models are trained by PyTorch on four Tesla V100 GPUs.
@@
Table 3: Training configurations for VIPTR.
Comparison with State-of-the-Art
@@
As indicated in Table 6, we compare VIPTR with prior studies on six English and one Chinese scene benchmarks. Even the fastest variant, VIPTRv2-T, achieves highly competitive accuracy, particularly among models without Language Models (LM). In English datasets, VIPTRv1-L outperforms LM-used models and SVTR-L, even achieving the best accuracy on IC15. Both VIPTRv1-L and VIPTRv2-B demonstrate overall accuracy comparable to recent studies Fang et al. (2021); Tang et al. (2021); Du et al. (2022); Lyu et al. (2022) while maintaining simplicity and faster execution.
@@
Turning to the Chinese Scene Dataset, the accuracy of 8 existing methods is provided by Du et al. (2022). Encouragingly, VIPTR performs notably well, with a 0.5% accuracy gain compared to SVTR-L, the best-performing model among those listed. Other VIPTR variants also exhibit advanced Chinese recognition accuracy. These results are attributed to VIPTR’s superior ability to perceive multi-grained character component features and its breakthroughs in characterizing Chinese words with rich stroke patterns.
4.1 Further Analysis
Ablation Experiments

### v5

@@ 3.4. Character and String Modeling
Existing studies mainly use feature sequences with height $1$ to represent image text. Each feature corresponds to a small local region. However, these segments are usually noisy, especially for irregular text, failing to describe the characters of an image well. In contrast, vision Transformer introduces two-dimensional feature representation, which not only enables more accurate discrimination between different characters and backgrounds but also fully exploits the potential contextual dependencies in patch sequence.
Figure 3. Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
[…]
Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
@@
More specifically, image text requires complementary features to be extracted from two perspectives. The first involves local visual features, representing internal associations and morphological features of each character. The second is contextual semantic relations, such as correlations between different characters or between text and non-text components. Therefore, we design two classes of self-attention blocks with different receptive fields to sense relevance. Mathematically speaking, a sequence of character components $B_{i-1}B_{i}$ in the previous stage, is first reshaped as a sequence of features. Upon inputting the self-attention block, Layer Normalization (LN) and Position Embedding (PE) are first applied. Subsequently, the self-attention mechanism is employed for dependency modeling. Finally, LN and Feed Forward Network (FFN) are successively applied for feature fusion, with the residual connection applied.
@@
Positional Embedding. Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back. Typical positional Embedding mechanisms include absolute positional encoding (APE) (Vaswani et al., 2017), relative positional encoding (RPE) (Liu et al., 2021), and CPE. APE and RPE are designed for a specific input size and are not friendly to varying input resolutions. CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions. Then the generated positional encoding will be added to the input feature. Locally-enhanced Positional Encoding (LePE) (Dong et al., 2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
@@
Local Attention. The goal of Local Attention is to encode character morphological features and establish correlations within character components, thereby simulating stroke features. Local Attention needs to consider the neighborhood of character components on the two-dimensional coordinate system. Drawing on the design ideas of MaSA and CSWin Self-Attention, we designed two efficient self-attention mechanisms that explicitly model local features on a two-dimensional coordinate system. MaSA is shown in Figure 3dLABEL:sub@subfig:dmsa. The spatial attenuation matrix of the image means that character features are highlighted globally and background features are weakened. CSWin Self-Attention is shown in Figure 3dLABEL:sub@subfig:cswin. This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel. For image text input, considering that the image resolution is larger than the Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing, and the length and width resolution needs to be guaranteed. Divisible by $SW$, we empirically set $SW$ for stages using CSWin Self-Attention to $1$ or $2$ by default.
@@
Global Attention. As shown in Figure 3dLABEL:sub@subfig:mhsa, Global Attention calculates the contextual dependencies between all character components. Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters. In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components. In addition, to save computational costs, we introduced OSRA, as Figure 3dLABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information. The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
3.5. Height Dimension Reduction
@@
To save computational costs and extract significant features layer by layer, we added the Height Dimension Reduction operation (HDR) after each self-attention block in every stage of SVIPTR, like classic CNN backbones (Borisyuk et al., 2018). Using the output of the features from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively. In the previous stage, we then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$. In the final stage, we first use AdaptiveAvgPool to compress the height dimension to $1$, followed by Fully Connected Layers, Hardswish Activation, and Dropout. By doing so, the character components are further compressed into feature sequences, where each element is represented by a feature of length $C_{3}$. Using AdaptiveAvgPool avoids applying convolution to embeddings with very small dimensions, such as height $2$, while being able to apply a variety of initial input heights.
The HDR block halves the height while maintaining a constant width. It not only reduces the computational cost but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension. We also increase the channel dimension $C_{i}$ to compensate for the loss of information.
3.6. Prediction
@@
3.7. Vision Permutation and Architecture Variants
There are various ways to permutate the global and Local Attention blocks in each stage. In previous researchs (Du et al., 2022), it was found that placing the Local Attention block before the Global Attention block is beneficial to both feature mining of individual symbols and to guiding the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model.
As shown in Figure 2, in the 2-3 stages of SVIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called SVIPTRv1 and SVIPTRv2. Symbolically, $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks. $[L_{i}][L_{j}//G_{s}][G_{t}]$ means in the stage 2-3, Local and Global Attention blocks are carried in parallel. The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA. Others are defined similarly. Effect of the two kinds of blocks Permutations can be seen in ablation experiments.
We design several hyperparameters in SVIPTR, including channel depth and number of heads per stage, and number of mixing blocks. By varying them, SVIPTR architectures with different numbers of parameters can be obtained. In summary, we constructed four typical architectures, namely SVIPTRv1-T (Tiny), SVIPTRv2-T, SVIPTRv1-L (Large), and SVIPTRv2-B (Base). Its detailed configuration is shown in Table 3.
Table 1. Ablation study on Attention Blocks’ permutations.
@@ 3.7. Vision Permutation and Architecture Variants
CRNN (None) 91.1 69.4 0.69
SVTR-T 93.5 74.8 0.29
$[L_{1}][L_{1}G_{1}][G_{1}]$ 94.4 80 0.269
$[L_{1}][L_{1}G_{2}][G_{1}]$ 94.4 80.2 0.258
$[L_{2}][L_{2}G_{1}][G_{1}]$ 94.3 80.1 0.282
$[L_{2}][L_{2}G_{2}][G_{1}]$ 94.6 80.4 0.27
$[L_{1}][L_{2}G_{1}][G_{1}]$ 94.1 79.9 0.27
$[L_{2}][L_{1}G_{1}][G_{1}]$ 94.2 80.1 0.27
$[L_{1}][L_{2}G_{2}][G_{1}]$ 94 80 0.258
$[L_{2}][L_{1}G_{2}][G_{1}]$ 94.1 79.8 0.258
$[L_{1}][L_{1}//G_{1}][G_{1}]$ 92.8 75.6 0.195
$[L_{1}][L_{1}//G_{2}][G_{1}]$ 92.7 75.6 0.192
$[L_{2}][L_{2}//G_{1}][G_{1}]$ 92.9 75.5 0.197
$[L_{2}][L_{2}//G_{2}][G_{1}]$ 93.1 76.1 0.194
$[L_{1}][L_{2}//G_{1}][G_{1}]$ 92.3 75 0.196
$[L_{2}][L_{1}//G_{1}][G_{1}]$ 92.5 74.8 0.196
$[L_{1}][L_{2}//G_{2}][G_{1}]$ 92.2 75.2 0.193
$[L_{2}][L_{1}//G_{2}][G_{1}]$ 92.5 75.4 0.193
4. Experiments and Discussion
4.1. Experiment Settings
@@
Our English STR models are trained on two synthetic datasets, MJSynth (MJ) (Jaderberg et al., 2014, 2016) and SynthText (ST) (Gupta et al., 2016) trainsets (72k+55k), along with real-world data from COCO-Text (Gomez et al., 2017). Testing is conducted on six public benchmarks: ICDAR 2013 (IC13) (Karatzas et al., 2013) testset with 857 regular images, Street View Text (SVT) (Wang et al., 2011) testset with 647 images from Google Street View, IIIT5K-3000 (Mishra et al., 2012) sourced from websites, ICDAR 2015 (IC15) (Karatzas et al., 2015) comprising 1811 images captured via Google Glasses, Street View Text Perspective (SVTP) (Phan et al., 2013) with 639 images also from Google Street View, and CUTE80 (CUTE), which focuses on curved text recognition, with 288 images.
For the Chinese recognition task, we employ the Chinese Scene and Document Dataset (Yu et al., 2021). The former includes a scene text dataset with 509,164 training, 63,645 validation, and 63,646 test images, featuring 5,880 unique characters. The document text dataset comprises 400,000 training, 50,000 validation, and 50,000 test images, with 4,865 distinct characters. We use the validation set to select the best model, which is subsequently evaluated on the test set.
Implementations. The recommended training and testing configurations for SVIPTR are listed in Table 2. When Training English models, data augmentation like stretch, perspective distortion, blur, and Gaussian noise, are randomly performed with probability $0.4$. Then, It was not used when training the Chinese model. All models are trained by PyTorch on four Tesla V100 GPUs.
Table 2. Training configurations for SVIPTR.
@@
Comparison with State-of-the-Art. As indicated in Table 4, we compare SVIPTR with prior studies on six English and one Chinese scene benchmarks. Even the fastest variant, SVIPTRv2-T, achieves highly competitive accuracy, particularly among models without Language Models (LM). In English datasets, SVIPTRv1-L outperforms LM-used models and SVTR-L, even achieving the best accuracy on IC15. Both SVIPTRv1-L and SVIPTRv2-B demonstrate overall accuracy comparable to recent studies (Fang et al., 2021; Tang et al., 2021; Lyu et al., 2022) while maintaining simplicity and faster execution.
@@
Turning to the Chinese Scene Dataset, the accuracy of 8 existing methods is provided by (Du et al., 2022). Encouragingly, SVIPTR performs notably well, with a 0.5% accuracy gain compared to SVTR-L, the best-performing model among those listed. Other SVIPTR variants also exhibit advanced Chinese recognition accuracy. These results are attributed to SVIPTR’s superior ability to perceive multi-grained character component features and its breakthroughs in characterizing Chinese words with rich stroke patterns.
Table 3. SVIPTR architecture variants (w/o counting the rectification module and CTC decoder).
SVIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 4.0 0.26
SVIPTRv1-L $[192,256,512,384]$ $[3,7,2,9]$ $[6,8,8,16]$ $[L_{2}][L_{2}G_{2}][G_{1}]$ 37.7 2.31
SVIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.2 0.19
SVIPTRv2-B $[128,256,384,256]$ $[3,6,6,9]$ $[4,8,8,12]$ $[L_{2}][L_{2}//G_{2}][G_{1}]$ 20.2 1.18
Table 4. Results on six English and Chinese scene benchmarks tested against existing methods. The model types from top to bottom are CTC-based, attention-based, LM-based, and using dedicated decoders. * means the lightweight version of the corresponding model. The speed is the average inference time on one Tesla V100 GPU with batch size 4 (to align with (Du et al., 2022)) for 3000 English image texts.
IC13 SVT IIIT5k IC15 SVTP CUTE
CRNN (Shi et al., 2016) 91.1 81.6 82.9 69.4 70.0 65.5 78.5 53.4 8.3 6.3
Rosetta (Borisyuk et al., 2018) 90.9 84.7 84.3 71.2 73.8 69.2 80.3 - 44.3 10.5
ViTSTR (Atienza, 2021) 93.2 87.7 88.4 78.5 81.8 81.3 85.6 - 85.5 11.2
SVTR-T (Du et al., 2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.0 4.5
SVTR-L (Du et al., 2022) 97.2 91.7 96.3 86.6 88.4 95.1 92.8 72.1 40.8 18.0
ASTER (Shi et al., 2018) 90.8 89.5 93.4 76.1 78.5 79.5 86.5 54.5 27.2 -
NRTR (Sheng et al., 2019) 94.7 88.3 86.5 74.1 79.4 88.2 84.0 - 31.7 160.0
SAR (Li et al., 2019) 91.0 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120.0
SATRN (Lee et al., 2020) 95.7 93.5 96.1 84.1 88.5 90.3 91.9 - - -
SRN (Yu et al., 2020) 95.5 91.5 94.8 82.7 85.1 87.8 90.4 60.1 54.7 25.4
ABINet (Fang et al., 2021) 97.4 93.5 96.2 86.0 89.3 89.2 92.6 - 36.7 51.3
VisionLAN (Wang et al., 2021a) 95.7 91.7 95.8 83.7 86.0 88.5 91.2 - 32.8 28.0
MATRN (Na et al., 2022) 95.8 94.9 96.7 82.9 90.5 94.1 92.3 - 44.2 29.6
MaskOCR* (Lyu et al., 2022) 97.7 93.7 95.4 86.6 89.0 87.5 92.5 68.6 36.0 24.1
MaskOCR (Lyu et al., 2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100.0 -
TrOCR (Li et al., 2023b) 98.3 93.2 91.0 84.0 91.0 89.6 90.2 - 558.0 20.8
MGP-STR-T (Wang et al., 2022) 96.5 93.2 96.3 86.2 89.4 90.6 92.7 - 5.4 10.6
MGP-STR(Fuse) (Wang et al., 2022) 97.3 94.7 96.4 87.2 91.0 90.2 93.3 - 21.0 12.0
SVTR-L+DCTC (Zhang et al., 2023) 97.4 93.7 96.9 87.3 88.5 92.3 93.3 73.9 40.8 18.0
SVIPTRv1-T (Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.9 4.2
SVIPTRv2-T 95.8 92.6 92.6 85.4 86.0 85.0 90.3 65.8 5.1 3.3
SVIPTRv1-L (Large) 97.3 94.5 95.1 88.4 89.6 88.8 92.9 72.6 39.5 16.9
SVIPTRv2-B (Base) 97.0 94.2 94.8 88.0 90.0 90.2 92.7 72.0 22.1 8.0
SVIPTRv1-L+BiLSTM+DCTC 98.0 94.8 96.1 88.8 90.4 91.2 93.6 74.2 45.8 22.0
4.2. Ablation Study